# Bidirectional Decoding: Improving Action Chunking via Closed-Loop Resampling

### v3

@@ D.3 Consistency-Reactivity Inequalities
$\left[P_{b}^{d}\left.P_{b}^{d}\mathbb{E}\left[\mathcal{L}(\pi_{(c+d,h)}(a_{t}|...,\hat{s}_{t-h-d-c:}^{t-h-c-1}=g_{t-h-d-c:}^{t-h-c-1}),\pi^{*})\Big{|}...,s_{t-h-d-c:}^{t-h-c-1}=g_{t-h-d-c:}^{t-h-c-1}\right]\right.\right.$
$+P_{\neq g_{t^{\prime}}}\left.P_{\hat{s}=s}^{b}\mathbb{E}\left[\mathcal{L}(\pi_{(c+d,h)}(a_{t}|...,\hat{s}_{t-h-d-c:}^{t-h-c-1}=s_{t-h-d-c:}^{t-h-c-1}),\pi^{*})\Big{|}...,s_{t-h-d-c:}^{t-h-c-1}\neq g_{t-h-d-c:}^{t-h-c-1}\right.\right]$
$+P_{b}^{d}\hat{P}_{\neq g_{t^{\prime}}}\left.\mathbb{E}\left[\mathcal{L}(\pi_{(c+d,h)}(a_{t}|...,s_{t-h-d-c:}^{t-h-c-1}\neq g_{t-h-d-c:}^{t-h-c-1}),\pi^{*})\Big{|}...,s_{t-h-d-c:}^{t-h-c-1}=g_{t-h-d-c:}^{t-h-c-1}\right.\right]$
$+{P}_{\neq g_{t^{\prime}}}P_{b}^{2d}\left.\mathbb{E}\left[\mathcal{L}(\pi_{(c+d,h)}(a_{t}|...,\hat{s}_{t-h-d-c:}^{t-h-c-1}=g_{t-h-d-c:}^{t-h-c-1}),\pi^{*})\Big{|}...,s_{t-h-d-c:}^{t-h-c-1}\neq g_{t-h-d-c:}^{t-h-c-1}\left.\right]\mid C,a_{t}\right]\mid C\right]$
$+{P}_{\neq g_{t^{\prime}}}P(\hat{s}_{t-h-d-c:}^{t-h-c-1}\neq g_{t-h-d-c:}^{t-h-c-1},{s}_{t-h-d-c:}^{t-h-c-1}\mid s_{t-h-c})$
@@
For the match terms, we use the fact that $P_{b}^{d}\leq 1$ and
$P(\hat{s}_{t-h-d-c:}^{t-h-c-1}={s}_{t-h-d-c:}^{t-h-c-1}|{s}_{t-h-c:}^{t-h})\leq 1$. For the mismatch terms, we use the definition of $\epsilon_{b}$ and Lemma 5. Then, we continue:
@@
$+P_{\neq g_{t^{\prime}}}\left[P_{b}^{d}\epsilon_{b}+\hat{P}_{\neq g_{t^{\prime}},s_{t^{\prime}}}\epsilon_{b}\right]+P_{b}^{d}\left[\hat{P}_{\neq g_{t^{\prime}}}\epsilon_{b}\right].$ (Simulation does not match reality)
@@
$\leq P_{\neq g_{t^{\prime}}}\left[P_{b}^{d}\epsilon_{b}+(1-P_{b}^{d})\epsilon_{b}\right]+P_{b}^{d}\left[\hat{P}_{\neq g_{t^{\prime}}}\epsilon_{b}\right]$
$=P_{\neq g_{t^{\prime}}}\epsilon_{b}+P_{b}^{d}\hat{P}_{\neq g_{t^{\prime}}}\epsilon_{b}$
$=P_{\neq g_{t^{\prime}}}\epsilon_{b}+\left.P_{b}^{d}\left[(1-P_{b}^{d})\right]\epsilon_{b}\right.$
$=(1-P_{b}^{d})\epsilon_{b}+\left.P_{b}^{d}\left[(1-P_{b}^{d})\right]\epsilon_{b}\right.$
$=\epsilon_{b}\cdot\left[1-P_{b}^{d}+P_{b}^{d}-P_{b}^{2d}\right]$
$=\epsilon_{b}\cdot\left[1-P_{b}^{2d}\right].$
@@
D.4 Discussion on Assumption 1
We assumed that $c+h<k$ so that the larger action chunk model can condition on more states from the distant past that are temporally correlated with $a_{t}$. In the case where $c+h\geq k$, the larger action chunk model will not get any advantage (thus, $\alpha=0$) and can only suffer from not having observed the recent past states. However, this is a very unlikely scenario because we expect human demonstrators to have a large memory horizon $k$.
D.5 Proof of Corollary 2 and Corollary 3
@@
Corollary 2 In a highly deterministic environment, if $a_{t}$ is temporally dependent on at least one state in $\{s_{t-h-c-d:t-h-c-1}\}$ and $\epsilon_{f}$ is finite, then
$\min_{\pi_{h+d}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h+d},\pi^{*})\mid C\right]<$ $\min_{\pi_{h}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h},\pi^{*})\mid C\right]$
@@
This follows from the upper bound of Proposition 1. We take $P_{f}\approx 1$ to get
$\min_{\pi_{h+d}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h+d},\pi^{*})\mid C\right]-\min_{\pi_{h}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h},\pi^{*})\mid C\right]\leq-\alpha_{b}+\epsilon_{f}(1-P_{f}^{2d})=-\alpha_{b}$
since $\epsilon_{f}$ is finite. Since $a_{t}$ is temporally dependent on at least one state in $\{s_{t-h-c-d:t-h-c-1}\}$, by Proposition 6, we have that $\alpha_{b}>0$. Therefore,
@@
Note that the lower bound becomes $\alpha_{f}-\epsilon_{b}(1-P_{b}^{2d})=-\epsilon_{b}(1-P_{b}^{2d})$ as $\alpha_{f}=0$ when $P_{f}=0$ (see Proposition 7). As such, we are guaranteed that both the upper and lower bounds are negative. ∎
Next, we prove Corollary 3. Note that $P_{b}$ has higher entropy since it is not conditioned on any action. In particular, we characterize a highly stochastic environment as one where $P_{b}$ is small due to the large amount of noise.
Corollary 3 In a highly stochastic environment, if temporal dependency decreases over time, i.e., $\alpha_{f}>\epsilon_{b}$, then
$\min_{\pi_{h+d}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h+d},\pi^{*})|C\right]>$ $\min_{\pi_{h}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h},\pi^{*})|C\right]$
@@
Starting from the lower bound of Proposition 1, since the environment is highly stochastic, we have that $P_{b}$ is small. We get:
$\min_{\pi_{h}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h}-\pi^{*})|C\right]\leq\min_{\pi_{h+d}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h+d}-\pi^{*})|C\right]-\alpha_{f}+\epsilon_{b}.$
By hypothesis, $\alpha_{f}>\epsilon_{b}$. Therefore, this reduces to:
@@
Note that the upper bound from Proposition 1 becomes $-\alpha_{b}+\epsilon_{f}$ as $P_{f}\approx 0$ in a highly stochastic environment. By Proposition 8, we have that $\epsilon_{f}-\alpha_{b}>\alpha_{f}-\alpha_{b}>\alpha_{f}-\epsilon_{b}>0$. Therefore, both the bounds are positive. ∎
@@
$-\epsilon_{b}(1-P_{b}^{2d})\leq-\alpha_{b}$
@@
$\alpha_{f}\leq\epsilon_{f}(1-P_{f}^{2d}).$
@@
$\quad+\min_{\pi_{h+d}}\Bigl{(}\mathbb{E}\bigl{[}P_{\text{correct inference}}\mathcal{L}\bigl{(}\pi_{h+d}(\text{correct inference}),\pi^{*}\bigr{)}\bigr{]}$
$\qquad+\mathbb{E}\bigl{[}P_{\text{incorrect inference}}\mathcal{L}\bigl{(}\pi_{h+d}(\text{incorrect inference}),\pi^{*}\bigr{)}\bigr{]}\Bigr{)}$
@@
$\quad+\min_{\pi_{h+d}}\Bigl{(}\mathbb{E}\bigl{[}P_{=g_{t^{\prime}}}P_{\text{correct inference}}\mathcal{L}\bigl{(}\pi_{h+d}(\text{correct inference}),\pi^{*}\bigr{)}\bigr{]}$
$\qquad+\mathbb{E}\bigl{[}P_{\neq g_{t^{\prime}}}P_{\text{correct inference}}\mathcal{L}\bigl{(}\pi_{h+d}(\text{correct inference}),\pi^{*}\bigr{)}\bigr{]}$
$\qquad+\mathbb{E}\bigl{[}P_{=g_{t^{\prime}}}P_{\text{incorrect inference}}\mathcal{L}\bigl{(}\pi_{h+d}(\text{incorrect inference}),\pi^{*}\bigr{)}\bigr{]}$
$\qquad+\mathbb{E}\bigl{[}P_{\neq g_{t^{\prime}}}P_{\text{incorrect inference}}\mathcal{L}\bigl{(}\pi_{h+d}(\text{incorrect inference}),\pi^{*}\bigr{)}\bigr{]}\Bigr{)}$
$\leq\epsilon_{b}(1-P_{b}^{2d}).$
D.6 Closed-Loop versus Open-Loop in Highly Stochastic and Near-Deterministic Environments
@@
In a highly stochastic environment, if temporal dependency decreases such that $\alpha_{f}>\epsilon_{b}$ at all time steps, then divergence between the closed-loop policy over the full trajectory is lower than that between the open-loop policy and the expert. In a near deterministic environment, if there is at least one time step $t$ such that $a_{t}$ depends on some state in $s_{t-d-c:t-c-1}$, then the divergence between the closed-loop policy over the full trajectory is greater than that between the open-loop policy and the expert.
@@
At any arbitrary time step $t$, the chunks of the two policies may be aligned in one of two ways:
Case 1: $\pi_{(c,1)}$ is executing $a_{t}$ as the first action in its action chunk and $\pi_{(c,1+d)}$ is also executing $a_{t}$ as the first action in its action chunk.
Case 2: $\pi_{(c,1)}$ is executing $a_{t}$ as the first action in its action chunk and $\pi_{(c,1+d)}$ is also executing $a_{t}$ as the $k$-th action, where $k\in(1,1+d]$ in its action chunk.
Using the Consistency-Reactivity Inequalities, in Case 1, both policies have equal divergence.
However, in case 2, using Corollary 3, we know that the closed-loop policy will outperform the open-loop one in the first setting of the statement and open-loop will outperform in the second. From this, we can conclude the divergence across the full trajectory. ∎

### v4

@@ E.3 Consistency-Reactivity Inequalities
$\left[\prod_{t^{\prime}\in\tau_{b}}P_{b}(t^{\prime})\left.\hat{P}_{b}(t^{\prime})\mathbb{E}\left[\mathcal{L}(\pi_{(c+d,h)}(a_{t}|...,\hat{s}_{t-h-d-c:}^{t-h-c-1}=g_{t-h-d-c:}^{t-h-c-1}),\pi^{*})\Big{|}...,s_{t-h-d-c:}^{t-h-c-1}=g_{t-h-d-c:}^{t-h-c-1}\right]\right.\right.$
$+P_{\neq g_{t^{\prime}}}\left.\hat{P}_{\hat{s}=s}^{b}\mathbb{E}\left[\mathcal{L}(\pi_{(c+d,h)}(a_{t}|...,\hat{s}_{t-h-d-c:}^{t-h-c-1}=s_{t-h-d-c:}^{t-h-c-1}),\pi^{*})\Big{|}...,s_{t-h-d-c:}^{t-h-c-1}\neq g_{t-h-d-c:}^{t-h-c-1}\right.\right]$
$+\prod_{t^{\prime}\in\tau_{b}}P_{b}(t^{\prime})\hat{P}_{\neq g_{t^{\prime}}}\left.\mathbb{E}\left[\mathcal{L}(\pi_{(c+d,h)}(a_{t}|...,s_{t-h-d-c:}^{t-h-c-1}\neq g_{t-h-d-c:}^{t-h-c-1}),\pi^{*})\Big{|}...,s_{t-h-d-c:}^{t-h-c-1}=g_{t-h-d-c:}^{t-h-c-1}\right.\right]$
$+{P}_{\neq g_{t^{\prime}}}\prod_{t^{\prime}\in\tau_{b}}\hat{P}_{b}(t^{\prime})\left.\mathbb{E}\left[\mathcal{L}(\pi_{(c+d,h)}(a_{t}|...,\hat{s}_{t-h-d-c:}^{t-h-c-1}=g_{t-h-d-c:}^{t-h-c-1}),\pi^{*})\Big{|}...,s_{t-h-d-c:}^{t-h-c-1}\neq g_{t-h-d-c:}^{t-h-c-1}\left.\right]\mid C,a_{t}\right]\mid C\right]$
$+{P}_{\neq g_{t^{\prime}}}\hat{P}_{\mathrm{train}}(\hat{s}_{t-h-d-c:}^{t-h-c-1}\neq g_{t-h-d-c:}^{t-h-c-1},{s}_{t-h-d-c:}^{t-h-c-1}\mid s_{t-h-c})$
@@
For the match terms, we use the fact that $\prod_{t^{\prime}\in\tau_{b}}\hat{P}_{b}(t^{\prime})\leq 1$ and
$\hat{P}_{\mathrm{train}}(\hat{s}_{t-h-d-c:}^{t-h-c-1}={s}_{t-h-d-c:}^{t-h-c-1}|{s}_{t-h-c:}^{t-h})\leq 1$. For the mismatch terms, we use the definition of $\epsilon_{b}$ and Lemma 5. Then, we continue:
@@
$+\prod_{t^{\prime}\in\tau_{b}}P_{b}(t^{\prime})(1-\hat{P}_{b}(t^{\prime}))\epsilon_{b}+\prod_{t^{\prime}\in\tau_{b}}(1-P_{b}(t^{\prime}))(\hat{P}_{b}(t^{\prime})+\hat{P}_{\neq g_{t^{\prime}},s_{t^{\prime}}})\epsilon_{b}.$ (Simulation does not match reality)
@@
$\prod_{t^{\prime}\in\tau_{b}}P_{b}(t^{\prime})(1-\hat{P}_{b}(t^{\prime}))\epsilon_{b}+\prod_{t^{\prime}\in\tau_{b}}(1-P_{b}(t^{\prime}))(\hat{P}_{b}(t^{\prime})+\hat{P}_{\neq g_{t^{\prime}},s_{t^{\prime}}})\epsilon_{b}$
$\leq\prod_{t^{\prime}\in\tau_{b}}P_{b}(t^{\prime})(1-\hat{P}_{b}(t^{\prime}))\epsilon_{b}+\prod_{t^{\prime}\in\tau_{b}}(1-P_{b}(t^{\prime}))(\hat{P}_{b}(t^{\prime})+(1-\hat{P}_{b}(t^{\prime})))\epsilon_{b}$
$=\epsilon_{b}\cdot\left[1-\prod_{t^{\prime}\in\tau_{b}}P_{b}(t^{\prime})\hat{P}_{b}(t^{\prime})\right]$
$=\epsilon_{b}\cdot\left[1-\prod_{t^{\prime}\in\tau_{b}}P_{b}(t^{\prime})(P_{b}(t^{\prime})+\delta_{b}(t^{\prime}))\right]$
@@
Corollary 2 (Consistency). Suppose, the train and test environments are the same and it is deterministic. Suppose $a_{t}$ is influenced by at least one state at time steps $\tau_{b}$ and let $\delta_{f}(t^{\prime})\approx 0$ for all $t^{\prime}\in\tau_{f}$. If the diversity in past strategies is not 0, and $\epsilon_{f}$ is finite, then
$\min_{\pi_{h+d}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h+d},\pi^{*})|C\right]<$ $\min_{\pi_{h}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h},\pi^{*})|C\right].$
@@
This follows from the upper bound of Proposition 1. Since the test environment is deterministic, we take $P_{f}(t^{\prime})\approx 1$ and, by hypothesis, $\delta_{f}(t^{\prime})\approx 0$ for all $t^{\prime}\in\tau_{f}$, menaing our learned policy’s implicit dynamics model is accurate. Then, we get
$\min_{\pi_{h+d}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h+d},\pi^{*})\mid C\right]-\min_{\pi_{h}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h},\pi^{*})\mid C\right]\leq-\alpha_{b}+\epsilon_{f}(1-\prod_{t^{\prime}\in\tau_{f}}P_{f}(t^{\prime})(P_{f}(t^{\prime})+\delta_{f}(t^{\prime})))=-\alpha_{b}$
since $\epsilon_{f}$ is finite. All that remains to show is that $\alpha_{b}$ is positive. Note that $a_{t}$ is temporally dependent on at least one state in $\{s_{t-h-c-d:t-h-c-1}\}$. Furthermore, since diversity in past strategies is not 0, the states in timesteps $\tau_{b}$ cannot be predicted with probability 1. Then, by Proposition 6, we have that $\alpha_{b}>0$. Therefore,
@@
Next, we prove Corollary 3.
Corollary 3 Suppose $P_{f}(t^{\prime})$ is small or both $P_{f}(t^{\prime})$ and $\left|\delta_{f}(t^{\prime})\right|$ are large for all $t^{\prime}\in\tau_{f}$. If temporal dependency decreases over time such that $\epsilon_{b}$ is small and $a_{t}$ is influenced by at least one state at time steps in $\tau_{f}$, then
$\min_{\pi_{h+d}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h+d},\pi^{*})|C\right]>$ $\min_{\pi_{h}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h},\pi^{*})|C\right].$
@@
Notice that since $\epsilon_{b}\approx 0$, the lower bound in Proposition 1 is $\alpha_{f}$. Next, we show that that $\alpha_{f}$ is positive. If $P_{f}(t^{\prime})$ is small, then we know by Proposition 7 that $\alpha_{f}$ is positive. On the other hand, if $P_{f}(t^{\prime})\approx 1$ but $\left|\delta_{f}(t^{\prime})\right|$ is large, then since probabilities are bounded, we require $\delta_{f}(t^{\prime})\approx-1$. In other words, the inference of the correct unobserved state is very unlikely causing $\alpha_{f}$ to be positive again. Therefore,
@@
$-\epsilon_{b}(1-\prod_{t^{\prime}\in\tau_{b}}P_{b}(t^{\prime})(P_{b}(t^{\prime})+\delta_{b}(t^{\prime})))\leq-\alpha_{b}$
@@
$\alpha_{f}\leq\epsilon_{f}(1-\prod_{t^{\prime}\in\tau_{f}}P_{f}(t^{\prime})(P_{f}(t^{\prime})+\delta_{f}(t^{\prime}))).$
@@
$\quad+\min_{\pi_{h+d}}\Bigl{(}\mathbb{E}\bigl{[}\mathcal{L}\bigl{(}\pi_{h+d}(\text{correct inference}),\pi^{*}\bigr{)}\bigr{]}\Bigr{)}$
$\qquad+\mathbb{E}\bigl{[}P_{\text{incorrect inference}}\mathcal{L}\bigl{(}\pi_{h+d}(\text{incorrect inference}),\pi^{*}\bigr{)}\bigr{]}$
@@
$\quad+\min_{\pi_{h+d}}\Bigl{(}\mathbb{E}\bigl{[}\mathcal{L}\bigl{(}\pi_{h+d}(\text{correct inference}),\pi^{*}\bigr{)}\bigr{]}\Bigr{)}$
$\qquad+\epsilon_{b}(1-\prod_{t^{\prime}\in\tau_{b}}P_{b}(t^{\prime})(P_{b}(t^{\prime})\delta_{b}(t^{\prime})))$
$=\epsilon_{b}(1-\prod_{t^{\prime}\in\tau_{b}}P_{b}(t^{\prime})(P_{b}(t^{\prime})+\delta_{b}(t^{\prime})))$
E.4 Closed-Loop versus Open-Loop in Highly Stochastic and Near-Deterministic Environments
@@
Suppose $P_{f}(t^{\prime})$ is small or both $P_{f}(t^{\prime})$ and $\left|\delta_{f}(t^{\prime})\right|$ are large for all $t^{\prime}\in\tau_{f}$. If temporal dependency decreases over time such that $\epsilon_{b}$ is small and $a_{t}$ is influenced by at least one state at time steps in $\tau_{f}$, then the expected divergence between the closed-loop policy over the full trajectory is lower than that between the open-loop policy and the expert.
Suppose, the train and test environments are the same and it is deterministic. Suppose $a_{t}$ is influenced by at least one state at time steps $\tau_{b}$ and let $\delta_{f}(t^{\prime})\approx 0$ for all $t^{\prime}\in\tau_{f}$. If the diversity in past strategies is not 0, and $\epsilon_{f}$ is finite, then the divergence between the closed-loop policy over the full trajectory is greater than that between the open-loop policy and the expert.
@@
At any arbitrary time step $t$, the chunks of the two policies can be aligned in one of two ways:
Case 1: $\pi_{(c,0)}$ is executing $a_{t}$ as the first action in its action chunk and $\pi_{(c,d)}$ is also executing $a_{t}$ as the first action in its action chunk.
Case 2: $\pi_{(c,0)}$ is executing $a_{t}$ as the first action in its action chunk and $\pi_{(c,d)}$ is executing $a_{t}$ as the $k$-th action, where $k\in(1,1+d]$ in its action chunk.
Using the Consistency-Reactivity Inequalities, in Case 1, both policies have equal divergence. However, in case 2, using Corollary 3, we know that the closed-loop policy will outperform the open-loop one in the first setting of the statement and open-loop will outperform in the second. From this we can conclude the divergence across the full trajectory. ∎